# MOTIF: Multi-strategy Optimization via Turn-based Interactive Framework

### v1

@@ D.1 Benchmark Datasets
For each problem domain, we synthesize training and test datasets, $\mathcal{D}\mathrm{train}$ and $\mathcal{D}\mathrm{test}$, following Ye et al. (2024), by sampling random instances under fixed settings. These datasets are used for fast evaluation during optimization.
@@ D.4 Evaluation Metrics
While the work of Dat, Doan, and Binh (2025) introduced two metrics for measuring code diversity, they are only applicable to a population of implementations that has not yet been clustered. In contrast, MOTIF naturally produces clustered implementations based on operator type. Accordingly, we adopt the following two metrics: novelty and silhouette score.
For each operator $o$, let $\mathcal{S}_{\pi}^{(o)}\subseteq\mathcal{S}_{\pi}$ denote the subset of implementations generated by applying $o$ to strategy $\pi$. We instantiate our embedding function $\mathcal{E}$ using the pretrained codet5p-110m-embedding (Wang et al. 2023), which maps each implementation into an $e$-dimensional vector (here $e=256$):
@@ D.5 Examples of Generated Outputs
Listing 1: Example implementation of generated construction strategy in DR
1# Final round optimized implementation for edge_score
2# Strategy ID: F1
3# Phase: Final round (system-aware)
4
5import numpy as np
6
7def edge_score(i: int, j: int, distances: np.ndarray) -> float:
8 """
9 Score for including edge (i,j) in greedy tour construction.
10 Higher score = more desirable edge for construction.
11 """
@@
13 weight_distance = 0.6
14 weight_connectivity = 0.4
15 connectivity_threshold = 25
16 n = distances.shape[0]
17
18 # Calculate distance score, shorter distances yield higher scores
19 distance_score = -distances[i, j] * weight_distance
20
21 # Count nearby unvisited cities within the threshold for potential future connections
22 connectivity_count = 0
23 for k in range(n):
24 if k != i and k != j:
25 if (distances[k, i] < connectivity_threshold) or (distances[k, j] < connectivity_threshold):
26 connectivity_count += 1
27
28 # Future edge bonus based on the number of nearby unvisited neighbors
29 future_edge_bonus = connectivity_count * weight_connectivity
@@
31 # Final score combining distance score with future connectivity potential
32 final_score = distance_score + future_edge_bonus
33 return np.clip(final_score, -1e6, 1e6) # Ensure score falls within reasonable bounds
Listing 2: Example implementation of generated deconstruction strategy in DR
1# Final round optimized implementation for city_badness
2# Strategy ID: F2
3# Phase: Final round (system-aware)
4
5import numpy as np
6
7def city_badness(tour_idx: int, tour: list[int], distances: np.ndarray) -> float:
8 n = len(tour)
9
10 # Hyperparameters
11 neighbor_weight = 2.0
12 disconnection_weight = 4.0
13 isolation_weight = 3.5
14 clustering_penalty_weight = 2.0
15 efficiency_weight = 1.2
16
17 city = tour[tour_idx]
18 neighbor_indices = [tour[(tour_idx + offset) % n] for offset in range(-2, 3) if offset != 0]
19
20 # Total distance to neighbors
21 total_neighbor_distance = np.sum([distances[city, neighbor] for neighbor in neighbor_indices])
22
23 # Clustering coefficient to penalize cities with too many connections
24 connections_count = np.sum(distances[city] < np.inf) - 1 # Exclude self
25 clustering_penalty = clustering_penalty_weight * (connections_count - 2)**2 if connections_count > 2 else 0
26
27 # Calculate isolation metrics
28 valid_neighbor_distances = [distances[city, neighbor] for neighbor in neighbor_indices if distances[city, neighbor] < np.inf]
29 avg_neighbor_distance = np.mean(valid_neighbor_distances) if valid_neighbor_distances else np.inf
30 overall_avg_distance = np.mean(distances[city][distances[city] < np.inf]) if np.sum(distances[city] < np.inf) > 1 else 0
31 isolation_penalty = isolation_weight * max(0, avg_neighbor_distance - overall_avg_distance)
32
33 # Efficiency metric that penalizes cities that contribute inefficient distances
34 route_efficiency = total_neighbor_distance / (len(valid_neighbor_distances) if valid_neighbor_distances else 1)
35 total_efficiency_penalty = efficiency_weight * max(0, route_efficiency - overall_avg_distance)
36
37 # Disconnection penalty
38 disconnection_penalty = (n - 1 - connections_count) * disconnection_weight
39
40 # Compute total badness score
41 total_badness = (total_neighbor_distance * neighbor_weight + isolation_penalty +
42 disconnection_penalty + clustering_penalty + total_efficiency_penalty)
43 return total_badness
Listing 3: Example implementation of generated repair strategy in DR
1# Final round optimized implementation for insert_position
2# Strategy ID: F3
3# Phase: Final round (system-aware)
4
5import numpy as np
6
7def insert_position(city: int, incomplete_tour: list[int], distances: np.ndarray) -> int:
8 best_position = 0
9 best_cost_increase = float(’inf’)
10 tour_length = len(incomplete_tour)
11
12 # Handle edge case where there is no tour
13 if tour_length == 0:
14 return 0
15
16 # Precompute distances for the city being inserted
17 city_distances = distances[city]
18
19 for position in range(tour_length + 1):
20 if position == 0:
21 cost_increase = city_distances[incomplete_tour[0]] # Before the first city
22 elif position == tour_length:
23 cost_increase = distances[incomplete_tour[-1], city] # At the end
24 else:
25 prev_city = incomplete_tour[position - 1]
26 next_city = incomplete_tour[position]
27 cost_increase = (distances[prev_city, city] + city_distances[next_city] - distances[prev_city, next_city])
28
29 # Update best position if current cost increase is better or prefer earlier position on ties
30 if cost_increase < best_cost_increase or (cost_increase == best_cost_increase and position < best_position):
31 best_cost_increase = cost_increase
32 best_position = position
33
34 # Early exit if we find a position with zero cost increase
35 if best_cost_increase == 0:
36 break
37
38 return best_position

### v2

@@ D.1 Benchmark Datasets
For each problem domain, we synthesize training and test datasets, $\mathcal{D}_{\text{train}}$ and $\mathcal{D}_{\text{test}}$, following ye2024reevo, by sampling random instances under fixed settings. These datasets are used for fast evaluation during optimization.
@@ D.4 Evaluation Metrics
While the work of dat2025hsevo introduced two metrics for measuring code diversity, they are only applicable to a population of implementations that has not yet been clustered. In contrast, MOTIF naturally produces clustered implementations based on operator type. Accordingly, we adopt the following two metrics: novelty and silhouette score.
For each operator $o$, let $\mathcal{S}_{\pi}^{(o)}\subseteq\mathcal{S}_{\pi}$ denote the subset of implementations generated by applying $o$ to strategy $\pi$. We instantiate our embedding function $\mathcal{E}$ using the pretrained codet5p-110m-embedding (wang2023codet5opencodelarge), which maps each implementation into an $e$-dimensional vector (here $e=256$):
@@ D.5 Examples of Generated Outputs
Below we present several representative strategy implementations generated by MOTIF during the multi-strategy optimization process for the Ant Colony Optimization (ACO) solver applied to the Travelling Salesman Problem (TSP). Rather than concentrating complexity into a single strategy component, MOTIF intentionally distributes the search across multiple routines, enabling the framework to capture improvements that originate from different parts of the solver. Consequently, the individual strategies produced by MOTIF may appear simpler compared to those generated by other frameworks, yet their collective contribution results in a much stronger overall algorithm.
Strategy 1: Initialization Strategy for Heuristic and Pheromone
d
1 n_cities = distances.shape[0]
2 distances_safe = distances + 1e-10 # Avoid division by zero
3
4 # Heuristic values based on normalized inverse distances to enhance desirability
5 heuristic = 1.0 / distances_safe
6 heuristic = (heuristic - np.min(heuristic)) / (np.max(heuristic) - np.min(heuristic) + 1e-10)
7
8 # Initialize pheromone with exponential decay based on distances
9 pheromone = np.exp(-distances_safe) # Exponential decay for pheromone initialization
10 pheromone_sum = pheromone.sum()
11 pheromone = pheromone / pheromone_sum if pheromone_sum > 0 else pheromone # Normalize
12 pheromone = np.clip(pheromone, 0.01, None) # Avoid zero threshold
13
14 return heuristic, pheromone
[…]
Strategy 2: Adaptive Transition Probability Computation
d
1 pheromone: np.ndarray,
2 heuristic: np.ndarray,
3 iteration: int,
4 n_iterations: int
5) -> np.ndarray:
6 # Dynamic hyperparameters
7 alpha = 1.0 + (2.0 * (iteration / n_iterations)) # Progressive increase
8 beta = 1.0 + (0.5 * (n_iterations - iteration) / n_iterations) # Decreasing influence
9 pheromone_decay = 0.95 # Pheromone decay factor
10 pheromone_min = 1e-10 # Minimum pheromone level to avoid zero influence
11
12 # Decay the pheromone values and ensure a minimum level
13 pheromone *= pheromone_decay
14 pheromone = np.maximum(pheromone, pheromone_min)
15
16 # Calculate unnormalized weights
17 weights = np.power(pheromone, alpha) * np.power(heuristic, beta)
18
19 # Normalize weights to ensure valid probabilities, avoiding division by zero
20 weights_sum = np.sum(weights, axis=1, keepdims=True)
21 weights_sum = np.where(weights_sum == 0, 1, weights_sum) # Default sum to one if zero
22 weights /= weights_sum # Normalize across rows to maintain proportions
23
24 return weights
[…]
Strategy 3: Pheromone Update with Iteration-Aware Reinforcement
d
1 pheromone: np.ndarray,
2 paths: np.ndarray,
3 costs: np.ndarray,
4 iteration: int,
5 n_iterations: int,
6) -> np.ndarray:
@@
8 min_pheromone = 0.01
9 evaporation_rate = 0.85
10 max_deposit = 50.0 # Capping maximum deposit level
11
12 # Evaporation process
13 pheromone *= evaporation_rate
14 pheromone = np.clip(pheromone, min_pheromone, None) # Avoid negative values
15
16 n_cities, n_ants = paths.shape
17 optimal_cost = np.min(costs) # Get the best cost to scale deposition
18
19 # Contribution and weighting
20 for ant in range(n_ants):
21 tour = paths[:, ant]
22 # Calculate pheromone deposit based on cost ratios and iterations
23 deposit = (optimal_cost / costs[ant]) * (1 + (iteration / (n_iterations + 1)))
24 deposit = np.clip(deposit, min_pheromone, max_deposit) # Ensure deposit caps
25 for i in range(n_cities):
26 c = tour[i]
27 n = tour[(i + 1) % n_cities]
28 pheromone[c, n] += deposit
29 pheromone[n, c] += deposit
@@
31 # Ensure pheromone levels do not fall below the minimum
32 pheromone = np.clip(pheromone, min_pheromone, None)
33 return pheromone
[…]
Appendix E Extended Discussions
E.1 Comparative Advantages
MOTIF departs from prior AHD frameworks by reframing solver design itself, emphasizing interactive search dynamics rather than isolated heuristic tuning. Its turn-based dual-agent structure introduces a peer-evaluation loop, while the two-phase workflow bridges localized improvements with system-level coherence. Together, these aspects give MOTIF several distinctive advantages:
  • Reformulating the problem. MOTIF shifts the goal from optimizing a single heuristic to jointly improving a set of interacting components, effectively introducing a new optimization problem: multi-strategy solver design.
  • Peer-driven refinement. Two LLM agents iteratively critique and outperform one another, creating a lightweight peer-review mechanism that yields more diverse revisions and stronger feedback than one-sided reflective prompting.
  • Local-to-global progression. The framework first encourages competitive, component-wise exploration and then transitions to cooperative, system-aware refinement, enabling MOTIF to escape local optima while still aligning all strategies into a coherent solver.
E.2 Limitations and Future Work
Limitations.
Although MOTIF is generally robust across domains, several factors may lead to temporary instability or reduced effectiveness. Most of these stem from the inherent variability of LLM-generated code and the dynamics of turn-based optimization, where small fluctuations in the agents’ behaviors can influence the search trajectory:
  • Non-deterministic code quality. LLM-generated implementations may occasionally fail to execute or degrade solver performance, especially when operators produce overly novel or structurally unconventional code.
  • Possible stalling in CMCTS. When operators repeatedly generate revisions that do not surpass the dynamic baseline, the competitive tree may stall temporarily, reducing effective search depth until diversity increases again.
  • Synergy mismatch in the global phase. In Phase 2, full-system updates can sometimes degrade performance if a locally beneficial change interacts poorly with other components, especially when strategy dependencies are subtle or highly non-linear.
Future Work.
Future extensions of MOTIF may explore more principled mechanisms for managing the expanding search space that arises when solvers are decomposed into many interacting components. This includes developing explicit state-management or version-control layers to track code variants efficiently, as well as methods for estimating the relative influence of each component so that search resources can be allocated adaptively to the most impactful routines. Another promising direction lies in learning or discovering better ways to factor a solver into modular sub-strategies—potentially allowing MOTIF to operate over more meaningful decompositions and reduce redundancy in the search structure.
E.3 Practical Applicability and Use Cases
MOTIF is not tied to any particular solver architecture, making it applicable well beyond the benchmark domains used in our experiments. Because the framework treats each strategy as an interchangeable module with a fixed signature, it can be plugged into a variety of classical, neural, or hybrid optimization pipelines. Several practical use cases naturally emerge:
  • Enhancing modular solvers. Many real-world optimization systems (e.g., routing engines, scheduling platforms, simulation–optimization loops) consist of distinct routines such as construction rules, update policies, or neighborhood operators. MOTIF can refine these modules without redesigning the surrounding pipeline.
  • Prototyping algorithmic ideas. Because the two agents rapidly explore alternative implementations, practitioners can use MOTIF as a tool for fast heuristic prototyping, obtaining interpretable code variants that reveal useful patterns even outside the automated search.
  • Automating domain adaptation. When a solver is deployed in a new setting—different instance distributions, constraints, or scaling regimes—MOTIF can generate strategy variants tailored to the new domain without requiring manual heuristic engineering.
E.4 Reproducibility Notes
The reproducibility of MOTIF is necessarily partial, as the framework depends on black-box LLMs whose sampling variability and internal updates introduce non-determinism into both code generation and search dynamics. While we provide full implementations, fixed prompts, strategy signatures, and evaluation settings, identical runs may diverge due to model randomness, operator exploration, or slight differences in API behavior. To support stable comparisons, we specify all hyperparameters, dataset generation procedures, and solver configurations used in our experiments; however, exact replication of every intermediate trajectory is not guaranteed.